a man buys 36 coins all of which are either 50 cents or 10 cents and the total value is 11.60 dollars how many stamps of each kind does he buy
Answer by mananth(16949)

x= number of 50 cent stamps
y= number of 10 cent stamps

1 x + 1 y = 36 .............1

50 x + 10 y = 1160 .............2
Eliminate y
multiply (1)by -10
Multiply (2) by 1
-10 x -10 y = -360
50 x + 10 y = 1160 (cents)
Add the two equations
40 x = 800
/ 40
x = 20
plug value of x in (1)
1 x + 1 y = 36
20 + y = 36
y = 36 -20
y = 16
y = 16 1
x= 20 number of 50 cent stamps
y= 16 number of 10 cent stamps
m.ananth@hotmail.ca